40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16960     Entry time: Tue Jun 28 22:27:11 2022
Author: Yehonathan 
Type: Update 
Category: BHD 
Subject: Measurement of input and output electronics noise 

{Yuta, Yehonathan}

For MICH noise budgeting we measure the input electronics noise which includes the AS55 RFPD, preamp, demod board, the whitening, and the AA filters, and the ADC noises. To do so we simply close the laser shutter and take the spectrum of C1:LSC-AS55_I_ERR_DQ and C1:LSC-AS55_Q_ERR_DQ shown in attachment 1.

Next, we measured the output electronics noise which includes the DAC, dewhitening and AI filters, and coil driver noises. We disabled the BS watchdog and went to 1X4 rack. We measured the spectrum of one of the lemo outputs on the BS coil driver module using an SR785. Attachment 2 shows the spectrum together with the SR785 dark noise.

Attachment 1: input_noise_spectrum.pdf  31 kB  Uploaded Tue Jun 28 23:58:19 2022  | Hide | Hide all
input_noise_spectrum.pdf
Attachment 2: output_noise_spectrum.pdf  21 kB  Uploaded Wed Jun 29 00:08:54 2022  | Hide | Hide all
output_noise_spectrum.pdf
ELOG V3.1.3-